katherine is 2 years older than christine. the sum of their ages is 16 how old is each girl

Answer by ikleyn(52803)

.

    Katherine + Christine = 16


       k     + (k - 2)   = 16

           2k            = 16 + 2 = 18

            k                     = 18/2 = 9.


ANSWER.  K is 9 years old.   C is 16-9 = 7 years old.

Solved.